Given f(x)=(ax)/(x^2+b), a≠0, b≠0. Find a and b such that f’(1)=0 and f’(0)=2.
Answer by edjones(8007)

Given f(x)=(ax)/(x^2+b), a≠0, b≠0. Find a and b such that f’(1)=0 and f’(0)=2.
.
when f(1)=0
a/(1+b)=0
a=(1+b)*0
a=0
b=any real number except -1 because division by zero is not allowed.
.
when f(0)=2
(a*0)/(0+b)=2
The equation cannot equal 2 it can only equal zero with any a or b.
.
Ed